OpenStudy (anonymous):

what is the solution to the following system y=1/2x+3 x=15-y

OpenStudy (istim):

Rearrange either the first equation to solve for x, or the 2nd equation to solve for y. Then sub the rearranged equation into the other, unchanged equation.

OpenStudy (blurbendy):

You know x = 15 - y, so we can use that in the first equation. y = 1/2(15 - y) + 3 y = 7.5 - 1/2y +3 y = 10.5 - 1/2y 1.5y = 10.5 y = 7 Now put y = 7 into the second equation. x = 15 - 7 x = 8 Put x = 8 and y = 7 into both equations and check your answer.

OpenStudy (istim):

A bit wordy there Blur. Try and let the student do their studying for themselves. They are potentially good enough.